# IMAGHarmony: Controllable Image Editing with Consistent Object Quantity and Layout

### v1

@@ 3.2 Harmony-Aware Attention Module
where $F_{v}^{\prime}\in\mathbb{R}^{B\times N\times D_{t}}$ and $W_{\text{proj}}\in\mathbb{R}^{D_{v}\times D_{t}}$ is a learnable matrix. Next, according to Eq. (2), we feed $F_{v}^{\prime}$ as the query and $F_{a}$ as both key and value into a multi-head cross-attention module to capture quantity-and-layout dependencies:
@@
where $F_{va}^{\prime}\in\mathbb{R}^{B\times N\times D_{v}}$ and $W_{\text{align}}\in\mathbb{R}^{D_{t}\times D_{v}}$ is a trainable projection. Furthermore, to flexibly control the influence of the injected harmony features and preserve the integrity of the original visual representations, we introduce a scaling hyperparameter $\alpha$. This parameter controls the strength of $F_{va}^{\prime}$ in the final representation $F_{h}\in\mathbb{R}^{B\times N\times D_{v}}$, as defined in Eq. (4).
$F_{h}=\alpha F_{va}^{\prime}+F_{v}.$ (4)
@@
The harmony feature $F_{h}$ encodes object quantity and layout, yielding a coherent and semantically consistent representation for image editing.
@@
Combined with SDXL. To inject harmony feature while preserving the original SDXL architecture, we adopt a dual-branch IP attention. One branch retains the frozen cross-attention module from SDXL, conditioned on a <Null Prompt> to avoid redundant textual interference. In parallel, the second branch incorporates the harmony feature $F_{h}\in\mathbb{R}^{B\times N\times D_{v}}$, which encodes object count and spatial layout. Then, we insert this IP attention module only at the “Down4” block of the UNet, as shown in Figure 2 (a). To maintain consistency with the weight of the pre-train IP-Adapter, the harmony feature $F_{h}$ is first projected through a frozen projection layer originally used for visual conditioning. This avoids introducing trainable parameters in the modality alignment stage and ensures compatibility with the pre-trained space. The projected harmony feature $F_{h}^{{}^{\prime}}$ is then fed into a learnable cross-attention block to interact with the query features. Let $F_{t}\in\mathbb{R}^{B\times N\times D_{t}}$ denote the textual feature, according to Eq. 5, the final attention output is computed.
@@
$Z_{\text{new}}=\text{Softmax}\left(\frac{QK_{t}^{T}}{\sqrt{d}}\right)V_{t}+\beta\cdot\text{Softmax}\left(\frac{QK_{h}^{T}}{\sqrt{d}}\right)V_{h},$ (5)
@@
where $Q=Z_{t}W_{q}$ is the query projected from the noisy latent $Z_{t}$. The text branch uses frozen projections $K_{t}=F_{t}W_{k}^{t}$, $V_{t}=F_{t}W_{v}^{t}$, while the harmony branch uses $K_{h}=F_{h}^{{}^{\prime}}W_{k}^{h}$, $V_{h}=F_{h}^{{}^{\prime}}W_{v}^{h}$ with newly introduced trainable weights. The hyperparameter $\beta$ balances the contribution of the harmony feature. This dual-branch design enables our method to inject structural semantics into the image editing process while preserving the generalization capabilities of the original SDXL model.
@@ 3.3 Preference-Guided Noise Selection Strategy
Motivation. Diffusion models are highly sensitive to initial noise [14, 63], especially in multi-object editing where object count and layout must be preserved. Even with the same prompt, different seeds often produce inconsistent results. To address this, we introduce a preference-guided noise selection (PNS) strategy that uses a frozen vision-language model to evaluate low-step generations and select the most semantically aligned seed, improving consistency and stability during inference.
@@
Seed Search. As shown in Figure 2 (a), we perform a seed search during the data preparation stage to pre-select structurally reliable initializations. Specifically, we uniformly sample 100 random seeds $\{s_{i}\}_{i=1}^{100}$, and for each seed $s_{i}$, we generate an image $I_{i}$ using SDXL [46] under a fixed textual prompt. To assess semantic fidelity, we employ CogVLM2 [19] as an evaluator by prompting it with instance-specific questions, such as “How many dogs are in the image?”. The predicted count is compared against the ground-truth, and a binary accuracy score $A_{i}\in\{0,1\}$ is assigned based on correctness. After computing $A_{i}$ for all seeds, we sort the results and select the top-$k$ seeds with the highest semantic alignment to form a candidate set $\mathcal{S}_{\text{top}}$. In practice, we set $k=5$, which balances efficiency and diversity. This seed search strategy ensures that inference begins from semantically aligned noise conditions, reducing layout distortion and object count errors in multi-object editing.
Seed Selection. As illustrated in Figure 2 (b), we utilize the candidate seed set $\mathcal{S}_{\text{top}}$ identified during the seed search phase. For each seed $s_{j}\in\mathcal{S}_{\text{top}}$, we initialize the diffusion model with the candidate noise and perform a shallow denoising process (10 steps) under the given editing instruction (e.g., “Cats”) to generate a candidate image $\hat{I}_{j}$. Each candidate image $\hat{I}_{j}$ is evaluated using CogVLM2 [19], which produces a semantic alignment score $R_{j}\in[0,1]$ reflecting its consistency with the editing instruction. We then identify the best candidate image with the highest alignment score and the best candidate noise. Finally, the best candidate noise is used to reinitialize the model, and a denoising process (30 steps) is performed to obtain the final edited image, which ensures output stability.
@@ 3.4 Training and Inference
Training. We optimize only the proposed HA module and IP attention of "Down4" block, while keeping the UNet of SDXL entirely frozen to retain its pre-trained generative capacity. Given the text features $F_{t}$ and the harmony features $F_{h}$, we adopt the standard MSE loss $L_{\text{MSE}}$ [12] defined as:
$L_{\text{MSE}}=\mathbb{E}_{x_{0},\epsilon,F_{t},F_{h},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},F_{t},F_{h},t)\right\|^{2},$ (6)
@@
Inference. During inference, we adopt the proposed PNS strategy to identify the best candidate. Building on the selected candidate, we further incorporate classifier-free guidance (CFG) [18] to enhance conditional generation and ensure consistency in object count and spatial layout. The final denoising prediction is computed as:
$\hat{\epsilon}_{\theta}(x_{t},F_{t},F_{h},t)=w\cdot\epsilon_{\theta}(x_{t},F_{t},F_{h},t)+(1-w)\cdot\epsilon_{\theta}(x_{t},t),$ (7)
@@
Figure 3: Examples from HarmonyBench. Each sample includes a source image and a validated count-object caption, verified by YOLO-World [4] and human review.
4 Experiments
@@
Datasets. To fill the gap in benchmarking quantity-and-layout consistent image editing (QL-Edit), we introduce a HarmonyBench dataset for evaluating a model’s ability to preserve object count and spatial layout during editing. The dataset contains 200 image-caption pairs, systematically organized by object count from 1 to 20, with 10 diverse layout examples per count level to ensure broad coverage of numerical and spatial variations. Images are sourced from real-world datasets (COCO [28], OpenImages [24]) and synthesized using SDXL [46], offering rich diversity in visual style, scene composition, and object arrangement. As shown in Figure 3, each caption (e.g., “Two Foxes” and “Twelve Dogs”) is verified through a two-stage pipeline: automatic object counting via YOLO-World [4], followed by manual inspection to ensure count accuracy and object clarity. During evaluation, each reference image is paired with 10 diverse editing prompts covering category changes, scene modifications, and style variations, resulting in a total of 2,000 test instructions.
@@
Metrics. We conducted a comprehensive evaluation of QL-Edit using both objective and subjective metrics. Objective metrics include object count accuracy (OCA) [34], which measures whether the number of generated objects matches the target count; average precision (AP) [6], which quantifies the spatial alignment between generated and reference layouts; and image reward (IR) [62], a learned score that assesses visual quality and instruction alignment. Subjective evaluation includes semantic correctness (SC) [68] and overall quality (OQ) [3], both based on human preference ratings. SC measures whether the generated image faithfully reflects the editing instruction, while OQ captures the overall perceptual quality and user satisfaction. Higher is better for all metrics.
Table 1: Quantitative comparisons on HarmonyBench. IMAGHarmony achieves the best performance.
Method Class Editing ($\uparrow$) Scene Editing ($\uparrow$) Style Editing ($\uparrow$)
OCA AP IR OCA AP IR OCA AP IR
MagicBrush [68] 70.3 69.2 0.437 91.3 90.5 0.104 90.8 76.7 0.216
InstructPix2Pix [3] 66.7 65.3 0.372 91.8 91.1 0.092 89.1 78.2 0.205
LEDITS [57] 45.2 43.1 0.463 77.2 71.9 -0.135 74.4 64.1 -0.253
LEDITS++ [2] 48.4 45.4 0.498 83.5 88.4 -0.187 77.9 63.9 -0.174
IC-Edit [69] 73.2 74.6 0.548 93.7 90.6 0.198 85.5 79.8 0.287
Show-o [61] 35.5 31.0 0.485 80.5 77.5 -0.134 30.6 43.6 0.124
SEED-X [11] 43.1 15.3 0.601 41.5 37.4 0.206 39.2 26.3 0.224
Emu-3 [59] 51.9 33.9 0.585 57.7 66.8 -0.041 45.1 31.2 0.268
IMAGHarmony (Ours) 80.7 75.7 0.693 94.3 91.8 0.217 91.4 90.2 0.315
@@
Implementation Details. All experiments are conducted on a single NVIDIA RTX 3090 GPU. Our configuration can be summarized as follows: (1) We initialize the denoising UNet with the pre-trained SDXL  [^1] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0, the IP attention with the pre-trained IP-Adapter  [^2] https://huggingface.co/h94/IP-Adapter. (3) The HA module adopts 8 attention heads by default. Input images are resized to $512\times 512$, and a 5% conditional dropout is applied to support classifier-free guidance (CFG). (4) We optimize the model using AdamW with a constant learning rate of $2.5\times 10^{-4}$ and train for 2k steps with a batch size of 1. During inference, we use the DDIM sampler with 30 denoising steps. Unless otherwise specified, the HA module scale, IP attention scale, and CFG scale are set to 1.0, 1.0, and 5.0, respectively.

### v2

@@ 3.3 Harmony Aware Module
where $F_{v}^{\prime}\in\mathbb{R}^{B\times N\times D_{t}}$ and $W_{\text{proj}}\in\mathbb{R}^{D_{v}\times D_{t}}$ is a learnable matrix. Next, according to Eq. (2), we feed $F_{v}^{\prime}$ as the query and $F_{a}$ as both key and value into a multi-head cross attention module to capture quantity-and-layout dependencies:
@@
where $F_{va}^{\prime}\in\mathbb{R}^{B\times N\times D_{v}}$ and $W_{\text{align}}\in\mathbb{R}^{D_{t}\times D_{v}}$ is a trainable projection. Furthermore, to flexibly control the influence of the injected perception features and preserve the integrity of the original visual representations, we introduce a scaling hyperparameter $\alpha$. This parameter controls the strength of $F_{va}^{\prime}$ in the final perception feature $F_{p}\in\mathbb{R}^{B\times N\times D_{v}}$, as defined in Eq. (4).
$F_{p}=\alpha F_{va}^{\prime}+F_{v}.$ (4)
@@
The perception feature $F_{p}$ encodes object quantity and layout, yielding a coherent and semantically consistent representation for image editing.
@@
To inject perception feature while preserving the original SDXL architecture, we adopt a dual-branch cross attention. One branch retains the frozen cross attention module from SDXL, conditioned on a <Null Prompt> to avoid redundant textual interference. In parallel, the second branch incorporates the perception feature $F_{p}\in\mathbb{R}^{B\times N\times D_{v}}$, which encodes object count and spatial layout. Then, we insert this cross attention module only at the "Down4" block of the UNet, as shown in Fig. 2 (a). Specifically, to use the weight of the pre-train IP-Adapter, the perception feature $F_{p}$ is first projected through a frozen projection layer originally used for visual conditioning. This avoids introducing trainable parameters in the modality alignment stage and ensures compatibility with the pre-trained space. The projected perception feature $F_{p}^{{}^{\prime}}$ is then fed into a learnable cross attention block to interact with the query features. Let $F_{t}\in\mathbb{R}^{B\times N\times D_{t}}$ denote the textual feature, according to Eq. 5, the final attention output is computed.
@@
$Z_{\text{new}}=\text{Softmax}\left(\frac{QK_{t}^{T}}{\sqrt{d}}\right)V_{t}+\beta\cdot\text{Softmax}\left(\frac{QK_{p}^{T}}{\sqrt{d}}\right)V_{p},$ (5)
@@
where $Q=Z_{t}W_{q}$ is the query projected from the noisy latent $Z_{t}$. The text branch uses frozen projections $K_{t}=F_{t}W_{k}^{t}$, $V_{t}=F_{t}W_{v}^{t}$, while the perception branch uses $K_{p}=F_{p}^{{}^{\prime}}W_{k}^{p}$, $V_{p}=F_{p}^{{}^{\prime}}W_{v}^{p}$ with newly introduced trainable weights. The hyperparameter $\beta$ balances the contribution of the perception feature. This dual-branch design enables our method to inject semantics of quantity and layout into the image editing process while preserving the generalization capabilities of the original SDXL model.
Discussion. In our training setup, the pretrained diffusion backbone remains fully frozen, and we fine-tune only the lightweight HA module; based on subsequent analysis, we inject it solely at the “Down4” block. This minimally intrusive adaptation avoids overfitting to the base distribution and catastrophic forgetting, while effectively preserving the backbone’s generalization and transferability. Meanwhile, systems represented by SDXL rely on CLIP-based encoders whose strength lies in global semantic alignment but whose modeling of instance-level structure (quantity and layout) is insufficient, often leading to repetition, omission, or layout drift in multi-object scenes. By introducing an auxiliary cross-modal path, HA provides object-level guidance for attention without modifying the backbone.
Figure 3: Examples of training data from HarmonyBench. Each sample includes a source image and a validated count-object caption, verified by YOLO-World [64] and human review.
@@ 3.4 Preference-Guided Noise Selection Strategy
Motivation. Diffusion models are highly sensitive to initial noise [30, 32], especially in multi-object editing where object count and layout must be preserved. Even with the same prompt, different seeds often produce inconsistent results. To address this, we present a preference-guided noise selection (PNS) strategy that uses a frozen vision-language model to evaluate low-step generations and select the most semantically aligned seed, improving consistency and stability during inference.
@@
Seed Search. As shown in Fig. 2 (a), we perform a seed search during the data preparation stage to pre-select structurally reliable initializations. Specifically, we uniformly sample N random seeds $\{s_{i}\}_{i=1}^{N}$, and for each seed $s_{i}$, we generate an image $I_{i}$ using SDXL [65] under a fixed textual prompt. To assess semantic fidelity, we employ vision–language model (VLM) as a judge by prompting it with instance-specific questions, such as “How many dogs are in the image?”. The predicted count is compared against the ground-truth, and a binary accuracy score $A_{i}\in\{0,1\}$ is assigned based on correctness. After computing $A_{i}$ for all seeds, we sort the results and select the top-$k$ seeds with the highest semantic alignment to form a candidate set $\mathcal{S}_{\text{top}}$. In practice, we set $N=100$ and $k=5$, which balances efficiency and diversity. This seed search strategy ensures that inference begins from semantically aligned noise conditions, reducing layout distortion and object count errors in multi-object editing.
Seed Selection. As illustrated in Fig. 2 (b), we utilize the candidate seed set $\mathcal{S}_{\text{top}}$ identified during the seed search phase. For each seed $s_{j}\in\mathcal{S}_{\text{top}}$, we initialize the diffusion model with the candidate noise and perform a shallow denoising process (10 steps) under the given editing instruction (e.g., “Cats”) to generate a candidate image $\hat{I}_{j}$. Each candidate image $\hat{I}_{j}$ is evaluated using CogVLM2 [66], which produces a semantic alignment score $R_{j}\in[0,1]$ reflecting its consistency with the editing instruction. We then identify the best candidate image with the highest alignment score and the best candidate noise. Finally, the best candidate noise is used to reinitialize the model, and a denoising process (30 steps) is performed to obtain the final edited image, which ensures output stability. Overall, during the seed search of the training stage, we quickly screen out candidate seeds that are roughly aligned in quantity through specific counting problems. In the subsequent seed selection of the inference stage, we use a more comprehensive semantic alignment score to select the best seed, ensuring that all aspects of the editing instructions are met. Discussion. The proposed PNS strategy targets a common yet under explored instability in diffusion based editing: different noise seeds can produce plausible images but with wrong counts or disordered layouts. Unlike InitNO [30] and Pick a Pic [60], which use vision language models to rank samples for global image–text alignment or aesthetics in generic generation, PNS focuses on structural constraints in multi-object editing. It formulates seed selection as a layout consistency aware ranking task by issuing layout sensitive queries (e.g., “How many dogs are in the image?” or “How many oranges are on the table?”) to a strong frozen VLM (e.g., CogVLM2 [66]) and scoring seeds by object count and spatial arrangement. Compared with inference time scaling for diffusion models (ITS DM) [67], which performs computationally intensive searches across multiple verifiers, subsequent experiments demonstrate that the PNS strategy is efficient and well-suited for layout-sensitive editing, delivering competitive results even with shallow denoising, such as 10 steps. Moreover, the PNS strategy also serves as a plug-and-play component for diffusion-based image editing frameworks.
TABLE III: Quantitative comparisons on proposed HarmonyBench. IMAGHarmony achieves the best performance.
[…]
SEED-X [58] 41.3 26.3 0.344 43.1 15.3 0.601 41.5 37.4 0.206 39.2 26.3 0.224
Show-o [59] 48.9 50.7 0.158 35.5 31.0 0.485 80.5 77.5 -0.134 30.6 43.6 0.124
Emu-3 [56] 51.6 44.0 0.271 51.9 33.9 0.585 57.7 66.8 -0.041 45.1 31.2 0.268
LEDITS [23] 65.6 59.7 0.025 45.2 43.1 0.463 77.2 71.9 -0.135 74.4 64.1 -0.253
LEDITS++ [25] 70.0 66.0 0.046 48.4 45.4 0.498 83.5 88.4 -0.187 77.9 63.9 -0.174
LoMOE [68] 72.3 67.9 0.059 66.4 63.8 -0.007 66.8 73.5 0.096 83.7 66.4 0.088
InstructPix2Pix [19] 82.5 78.2 0.223 66.7 65.3 0.372 91.8 91.1 0.092 89.1 78.2 0.205
MagicBrush [18] 84.1 78.8 0.252 70.3 69.2 0.437 91.3 90.5 0.104 90.8 76.7 0.216
IC-Edit [26] 84.1 81.7 0.344 73.2 74.6 0.548 93.7 90.6 0.198 85.5 79.8 0.287
IMAGHarmony (SD1.5) 88.3 85.5 0.401 79.6 75.6 0.686 94.1 91.4 0.204 91.2 89.4 0.312
IMAGHarmony (SDXL) 88.8 85.9 0.408 80.7 75.7 0.693 94.3 91.8 0.217 91.4 90.2 0.315
@@ 3.5 Training and Inference
Training. We optimize only the proposed HA module, while keeping the UNet of SDXL entirely frozen to retain its pre-trained generative capacity. Given the text features $F_{t}$ and the perception features $F_{p}$, we adopt the standard MSE loss $L_{\text{MSE}}$ [69] defined as:
$L_{\text{MSE}}=\mathbb{E}_{x_{0},\epsilon,F_{t},F_{p},t}\left\|\epsilon-\epsilon_{\theta}(x_{t},F_{t},F_{p},t)\right\|^{2},$ (6)
@@
Inference. During inference, we incorporate classifier-free guidance (CFG) [70] to enhance conditional generation and ensure consistency in object count and spatial layout. The final denoising prediction is computed as:
$\hat{\epsilon}_{\theta}(x_{t},F_{t},F_{p},t)=w\cdot\epsilon_{\theta}(x_{t},F_{t},F_{p},t)+(1-w)\cdot\epsilon_{\theta}(x_{t},t),$ (7)
@@
TABLE IV: Quantitative comparisons on proposed HarmonyBench according to the number of the object.
[…]
Show-o [59] 64.6 79.6 0.262 55.4 65.2 0.194 47.8 37.1 0.163 27.6 20.9 0.014
SEED-X [58] 68.3 53.1 0.407 47.7 29.8 0.401 26.2 12.7 0.322 22.9 9.7 0.245
Emu-3 [56] 74.8 61.2 0.305 61.1 47.2 0.302 42.6 37.5 0.285 27.7 29.9 0.214
LEDITS [23] 77.8 69.2 0.205 67.3 59.3 0.005 64.7 57.3 -0.047 52.6 52.9 -0.063
LEDITS++ [25] 78.2 70.8 0.199 71.9 68.4 -0.004 65.7 65.2 -0.005 63.9 59.2 -0.073
LoMOE [68] 88.6 80.6 0.236 72.1 74.4 0.203 68.0 63.7 -0.004 60.5 52.9 -0.055
InstructPix2Pix [19] 88.2 88.1 0.322 87.5 83.0 0.293 79.4 72.9 0.183 74.9 69.1 0.091
MagicBrush [18] 89.1 87.8 0.305 87.4 83.7 0.296 81.7 72.5 0.254 78.3 71.2 0.154
IC-Edit [26] 89.3 88.1 0.402 85.2 84.8 0.383 83.2 79.4 0.298 78.8 74.3 0.294
IMAGHarmony (SD1.5) 89.5 88.3 0.414 87.9 84.6 0.402 87.6 84.9 0.409 88.3 84.1 0.408
IMAGHarmony (SDXL) 89.9 88.5 0.419 88.4 85.2 0.404 87.7 84.8 0.411 89.2 85.1 0.399
4 Experimental
4.1 Implementation Details
@@
Datasets. To fill the gap in benchmarking quantity-and-layout consistent image editing (QL-Edit), we devise a HarmonyBench dataset for training and evaluating a model’s ability to preserve object count and spatial layout during editing. The training set consists of 0.2K image–caption pairs uniformly distributed over object counts: for counts one through nine, we include 20 distinct layout instances per count, and an additional 20 instances for counts greater than nine, ensuring broad coverage of numerical and spatial variations. Importantly, none of the pairs contain structural annotations or explicit multi-object labels; that is, the dataset provides no bounding boxes, masks, or layout information. Similarly, the evaluation set comprises 2K instruction–image pairs spanning class changes, scene modifications, and style variations. Noted that the training set and the test set do not overlap. The data construction process of HarmonyBench is as follows: Images are sourced from real-world datasets (COCO [71], OpenImages [72]) and synthesized using SDXL [65], offering rich diversity in visual style, scene composition, and object arrangement. As shown in Fig. 3, each caption (e.g., “Two Foxes” and “Twelve Dogs”) is verified through a two-stage pipeline: automatic object counting via YOLO-World [64], followed by manual inspection to ensure count accuracy and object clarity. To further validate the generalizability of our approach, we also conducted experiments on MagicBrush [18], a standard benchmark dataset for image editing.
@@
Metrics. We evaluate QL-Edit with a combination of objective and subjective metrics. On the objective side, we report: (i) object count accuracy (OCA) [73], which checks whether the number of generated objects matches the target count; (ii) average precision (AP) [74], which quantifies the spatial alignment between the generated and reference layouts; and (iii) image reward (IR) [75], a learned score assessing visual quality and instruction alignment. OCA and AP are computed using YOLO-World for counting and detection. Thus, OCA constrains object count, AP measures spatial matching, and IR captures holistic image quality; together they provide a complementary, comprehensive quantitative evaluation of QL-Edit. For subjective assessment, we adopt semantic correctness (SC) [18], whether the image faithfully reflects the editing instruction, and overall quality (OQ) [19], which reflects perceptual quality and user preference. All metrics are “higher is better.”
@@
Hyperparameters. All experiments are conducted on a single NVIDIA RTX 3090 GPU. Our configuration can be summarized as follows: (1) We initialize the denoising UNet with the pre-trained SDXL  [^1] https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0, the cross attention with the pre-trained IP-Adapter  [^2] https://huggingface.co/h94/IP-Adapter. (3) The HA module adopts 8 attention heads by default. Input images are resized to $512\times 512$, and a 5% conditional dropout is applied to support classifier-free guidance (CFG). (4) We optimize the model using AdamW with a constant learning rate of $2.5\times 10^{-4}$ and train for 2k steps with a batch size of 1. During inference, we default use the DDIM sampler with 30 denoising steps. Unless otherwise specified, the HA module scale, IP attention scale, and CFG scale are set to 1.0, 1.0, and 5.0, respectively.